The speed of light in a vacuum is 3.00 x 108 m/s. Calculate the speed of light in diamond (n=2.42).
1
Expert's answer
2020-12-21T11:00:32-0500

We can find the speed of light in diamond from the Snell's Law:


"\dfrac{v_2}{v_1}=\dfrac{n_1}{n_2},"

here, "v_1=3\cdot10^8\ \dfrac{m}{s}" is the speed of light in the vacuum, "v_2" is the speed of light in diamond, "n_1=1.0" is the index of refraction of the vacuum and "n_2=2.42" is the index of refraction of diamond.

Then, from this formula we can find the speed of light in diamond:


"v_2=\dfrac{v_1n_1}{n_2}=\dfrac{3\cdot10^8\ \dfrac{m}{s}\cdot 1.0}{2.42}=1.24\cdot10^8 \ \dfrac{m}{s}."

Answer:

"v_2=1.24\cdot10^8 \ \dfrac{m}{s}."